OpenStudy (anonymous):

if x+sqrt(x)=1 what is the value of x^2+2x.sqrt(x)

OpenStudy (anonymous):

\[x+\sqrt{x}=1\] \[\sqrt{x}=1-x\]x^2-3x+1=0 \[x=1-2x+x^2\] \[x^2-3x+1=0\] \[x=\frac{3-\sqrt{5}}{2}\] and \[1=(x+\sqrt{x})^2=x^2+2x\sqrt{x}+x\]so \[x^2+2\sqrt{x}=1-\frac{3-\sqrt{5}}{2}=\frac{2-3+\sqrt{5}}{2}=\frac{-1+\sqrt{5}}{2}\]

OpenStudy (anonymous):

where did you get x^2-3x+1=0 from? Oo

OpenStudy (anonymous):

\[\sqrt{x}=1-x\] square both sides get \[x=(1-x)^2=1-2x+x^2\] set equal to zero to solve the quadratic get \[x^2-3x+1=0\] unless i goofed up

OpenStudy (anonymous):

no, you're right, i just didn't pay attention :P

OpenStudy (anonymous):

you really in brazil?!

OpenStudy (anonymous):

haha yeah lol

OpenStudy (anonymous):

wow.

OpenStudy (anonymous):

near the ocean?

OpenStudy (anonymous):

10 min :P

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

so if i sometimes say something wrong in my questions, it's becaus i have to translate them all :)

OpenStudy (anonymous):

from portuguese?

OpenStudy (anonymous):

yup. all of them :)

OpenStudy (anonymous):

i am impressed

OpenStudy (anonymous):

be

OpenStudy (anonymous):

why not? i can barely speak one language, and portuguese is way different than english

OpenStudy (anonymous):

yeah but i lived in england for a few years when i was little, so i still remember a few things. The hard part is to know scientific words, but MIT open course helps me with that lol

OpenStudy (anonymous):

i bet brazil in nicer than england

OpenStudy (anonymous):

no way, i'm doing everything to go back. where are you from?

OpenStudy (anonymous):

philadelphia

OpenStudy (anonymous):

going back when you graduate? or sooner?

OpenStudy (anonymous):

sooner, my dad works in reasearch and he knows a lot of people at some universities in england so i might try to get a transfer or something when he goes there to do his Post-doc

OpenStudy (anonymous):

good luck! hope you get there soon.

OpenStudy (anonymous):

thank you :)